OpenStudy (anonymous):

Problem: The rectangular playing area of a professional football field is 100 yards long. The perimeter of the playing area is 920 feet. The formula for the perimeter of a rectangle is P=2L+2W P=920 L=100 , (here are two ways to solve this problem. You can solve the formula for the unknown variable, then substitute. Or, you can substitute, then solve the resulting equation)

OpenStudy (anonymous):

HELP PLEASE

OpenStudy (anonymous):

L=100 yards=100*3=300 feet Let width =x feet 920=2*300+2x 920-600=2x 2x=320 x=320/2=160 feet width=160 feet